How do you write the residue at z=∞ where f(z) = 1/z^3+z^5?


A)-2


B)-1


C)1


D)0


1
Expert's answer
2022-02-14T16:51:06-0500

To calculate the residue at infinity we have the following relation :

"Res_\infty f(z) = -Res_0 g(\xi)" with "g(\xi)=f(1/\xi)".

Therefore, we consider the residue of "\xi^3+1/\xi^5" at zero.

We can calculate it directly (for example, we know that the residue is the coefficient before "1/\xi" in the Laurent series) and we find that "Res_0 g(\xi)=0" and therefore "Res_\infty f(z)=0" , the correct answer is D.

A faster way to obtain this result could be using the fact that the sum of residues of a meromorphic function (counting the residue at the infinity) is zero. The only residue of "f" in "\mathbb{C}" is the residue at zero and it is null.